The Bookstall Inc. is a specialty bookstore concentrating on used books sold via the Internet. Paperbacks are $1.85 each, and hardcover books are $4.85. Of the 100 books sold last Tuesday morning, 85 were paperback and the rest were hardcover.

What was the weighted mean price of a book? (Round your answer to 2 decimal places.)

1. Weighted Mean $________
Answer #1

Since, as we know that Weighted mean.

Weighted mean is nothing but its is mean of sum of products of it's means and number of observations i.e.

ar_x = rac{N_p * W_p + N_h * W_h }{N_p + N_h}

Here, N_p = Number of paperback ,

N_h = Number of hardcover,

W_p = Mean of Cost of Paperback,

W_h = Mean of Cost of hardcover.

Given, Total number of paperback = 85

and total number of hardcover books = 100 - 85 = 15

Cost of each paperback = $1.85

impliesCost of 85 paperbacks = 85 * $1.85 = $157.25

Therefore ,Mean of cost of 85 paperbacks = 157.25 85 = $1.85

Now, Cost of each hardcover = $4.85

impliesCost of 15 hardcovers = 15 * $4.85 = $72.75

Therefore, Mean of cost of 15 hardcovers = 72.75 15 = $4.85

implies Using formula for weighted mean = 85 1.85 15 4.85 85 15

implies Weighted Mean = rac{157.25 + 72.75}{100} =  230 100 = 2.30. (upto 2 decimal places)